From the tree Let's say ? = (?, ?) is undirected graph, with self-loops or no parallel edges. And Let's say |?| = ? and |?| = ?.

Prove this thing by induction which 2? <= n2 − ? for all ? >= 1.(Its about the tree.)
Answer #1

Ans:-

Base case:

n= 2, m= 1

Because we can connect two vertices with only one edge.

This proves

2*m<= n^2 - n

2*1<= 2^2 - 2

=> 2<= 2

The base case holds true.

inductive case:

for n= 3, we can have m= 2 for fulfilling the condition of a simple graph

which means

2*2 <= 3^2 - 3

=> 4<= 9 - 3

=> 4<= 6

Similarly for greater value of m and n the inequality

2*m<= n^2 - n will hold true.